# Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v2

@@ 4.2 Dual-branch multi-modal joint training
Since information from different modalities is exchanged and learned through CICO and injected into the trainable visual backbone, we utilize the last-epoch visual branch for inference.
Output: Predicted label $\hat{y}_{i}$.
/* Stage 1: TES Training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
            $\boldsymbol{z}_{i}^{v}=E_{i}(\boldsymbol{x}_{i})$ // visual embedding
            $\boldsymbol{t}_{i}=l(\boldsymbol{z}_{i}^{v}))$ // pseudo text tokens
            $\mathcal{L}_{align}\leftarrow$Eqs. (5) and (6)
            $\mathcal{L}_{distill\leftarrow}$q. (7)
            $\mathcal{L}_{TES}\leftarrow$Eq. (8)
            Back-propagation and optimize $\theta_{\text{t}}$.
until reaching max epochs;
/* Stage 2: Dual-branch training */
repeat
       for  $(\boldsymbol{x}_{i},\boldsymbol{y}_{i})\in\text{each batch}$ do
             /* Visual-branch */
            Compute $\mathcal{L}_{ucon}^{v}$ and $\mathcal{L}_{scon}^{v}$ by replacing $\boldsymbol{h}$ in Eqs.(1) and (2) with $\boldsymbol{h}^{v}$
            $\mathcal{L}_{\text{rep}}^{v}\leftarrow$Eq. (9)
            Compute $\mathcal{L}_{\text{cls}}^{v}$ by replacing $\boldsymbol{p}$ in Eqs.(3) and (4) with $\boldsymbol{p}^{v}$
            $\mathcal{L}_{\text{db}}^{v}\leftarrow$Eq. (10)
            /* text-branch */
             $\boldsymbol{\hat{z}}_{i}^{t}=E_{t}(l(E_{i}(\boldsymbol{x}_{i})))$
            $\mathcal{L}_{\text{db}}^{t}\leftarrow$Eq. (11)
            Compute the multi-modal mean entropy regularization $H_{mm}$
            /* CICO */
             Calculate the visual and text anchors $\mathcal{P}_{v},\mathcal{P}_{t}$
            Compute the instance relationships by Eq. (12)
            $\mathcal{L}_{\text{CICO}}\leftarrow$Eq. (13)
            $\mathcal{L}_{\text{Dual}}\leftarrow$Eq. (14)
until reaching max epochs;
return $\hat{y}_{i}=\eta(f_{v}(\boldsymbol{x}_{i}))$.
Algorithm 1 Pseudocode for  GET.
@@ 5.1 Experimental setup
Datasets.  We evaluate our method on multiple benchmarks, including three image classification generic datasets (i.e., CIFAR 10/100 [31] and ImageNet-100 [32]), three fine-grained datasets from Semantic Shift Benchmark [33] (i.e., CUB [34], Stanford Cars [35] and FGVC-Aircraft [36]), and three challenging datasets (i.e., Herbarium 19 [37], ImageNet-R [38] and ImageNet-1K [32]). Notice that we are the first to introduce ImageNet-R into the GCD task, which contains various renditions of 200 ImageNet classes, thus challenging the GCD’s assumption that the data comes from the same domain. For ImageNet-R, we subsample the first 100 classes as old classes, leaving the rest as new classes; the labeled dataset $\mathcal{D}_{l}$ consists of half of the old class samples, while the other half and all the new class samples are used to construct unlabelled dataset $\mathcal{D}_{u}$. As for other benchmarks, we follow the previous [3, 4] to sample $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$. The details of the datasets we evaluate on are shown in Table I.
TABLE I: The details of the datasets we evaluate on.
@@
Evaluation metrics.  Following standard evaluation protocol (V2) in [3], we evaluate the performance with clustering accuracy (ACC), calculated as:
$\mathcal{ACC}=\frac{1}{|\mathcal{D}_{u}|}\sum_{i=1}^{|\mathcal{D}_{u}|}\mathbbm{1}{[y_{i}=perm(\hat{y}_{i})]}\,,$ (15)
where $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted label, respectively. $perm$ is the optimal permutation to obtain the highest ACC and can be computed with the Hungarian algorithm [39].
Implementation details.  We use a ViT-B/16 [16] pre-trained with CLIP [9] as the backbone of the image and text encoder. In the first stage, we train a fully connected layer to transfer image embeddings into pseudo tokens. In the second stage, we only fine-tune the last transformer block of the image encoder for all datasets and remove the projector resulting in features with a dimension of 768. The exception is ImageNet-1K, we remain and fine-tune the last projection layer, which avoids gradient explosion and improves results with lower computational cost, resulting in features with a dimension of 512. We use a single linear layer to turn pseudo text embeddings generated by TES into learnable embeddings while changing their dimensions (512 to 768) to match those of the visual features. The batch size is fixed to 128 for training and 256 for testing. Training is done with an SGD optimizer and an initial learning rate of 0.1 decayed by a cosine annealing rule. We train for 200 epochs on each dataset in both two stages. In the first stage, we set the number of pseudo text tokens to 7. The balance coefficient $\lambda$ is set to 0.35 as [3], while $\lambda_{c}$ is set to 2 and 1, respectively. The temperature value $\tau_{a}$ is set to 0.01 while other temperature values $\tau_{c}$, $\tau_{s}$ and $\tau_{t}$ are as same as [4]. All experiments are conducted with 4 NVIDIA GeForce RTX 3090 GPUs.
5.2 Comparison with state of the arts
In this section, we compare the proposed GET with the several state-of-the-art methods, including: $k$-means [40], RS+ [12], UNO+ [14], ORCA [41], PromptCAL [5], DCCL [42], GPC [43], GCD [3] and SimGCD [4]. GCD [3] and SimGCD [4] provide paradigms for non-parametric and parametric classification, thus we reproduce them by replacing their backbone with CLIP for a fair comparison, denoted as GCD-CLIP and SimGCD-CLIP. The results presented in the paper are the averages of three independent repeated runs. We provide the performance standard deviation of main results in the Appendix.
Evaluation on fine-grained datasets.  As shown in Tab. II, our method achieves consistently remarkable success on all three fine-grained datasets. Specifically, we surpass SimGCD-CLIP by 5.3%, 8.5%, and 4.6% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. In fine-grained datasets, the visual conceptions of distinct classes exhibit high similarity, making it challenging for classification based solely on visual information. However, their text information can provide additional discriminative information. Consequently, our GET  significantly enhances classification accuracy through the reciprocal augmentation of text and visual information flow.
TABLE II: Results (%) on fine-grained and generic datasets. The best results are highlighted in bold
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
TABLE III: Results (%) on more challenging datasets. The best results are highlighted in bold
Herbarium 19 ImageNet-1K ImageNet-R
Method All Old New All Old New All Old New
$k$-means [40] 13.0 12.2 13.4 - - - - - -
RS+ [12] 27.9 55.8 12.8 - - - - - -
UNO+ [14] 28.3 53.7 14.7 - - - - - -
ORCA [41] 20.9 30.9 15.5 - - - - - -
GCD [3] 35.4 51.0 27.0 52.5 72.5 42.2 32.5 58.0 18.9
SimGCD [4] 44.0 58.0 36.4 57.1 77.3 46.9 29.5 48.6 19.4
GCD-CLIP 37.3 51.9 29.5 55.0 65.0 50.0 44.3 79.0 25.8
SimGCD-CLIP 48.9 64.7 40.3 61.0 73.1 54.9 54.9 72.8 45.3
GET (Ours) 49.7 64.5 41.7 62.4 74.0 56.6 58.1 78.8 47.0
Evaluation on generic datasets.  In Table II, we also present the performances for three generic datasets. Due to the low resolution of the CIFAR dataset and model biases (CLIP itself performs poorly on CIFAR100, with a zero-shot performance of 68.7), the results for novel classes are inferior compared to the DINO backbone. However, despite the inherent limitations in the discriminative capability of CLIP itself, our method still achieves an improvement of 0.4% on ‘Old’ classes of CIFAR10 and 2.2% on ‘New’ classes of CIFAR100, compared to SimGCD-CLIP. For ImageNet-100, SimGCD-CLIP has achieved an exceptionally saturated result of 90.8% on ‘All’ classes, further advancements pose considerable challenges. However, leveraging the additional modality information, GET elevates the performance ceiling to an impressive 91.7%.
Evaluation on challenging datasets.  We further evaluate on three challenging datasets. As shown in Tab. III, GET outperforms all other state-of-the-art methods for both ‘All’ and ‘New’ classes on Herb19 and ImageNet-1K datasets. In particular, our method achieves a notable improvement of 1.4% and 1.7% on ‘New’ classes of Herb19 and ImageNet-1K, respectively. Furthermore, the suboptimal performance of GCD and SimGCD with the DINO backbone on the ImageNet-R dataset highlights the difficulty of DINO in discovering new categories with multiple domains. Despite multiple domains for images of the same category, their textual information remains consistent. Our method effectively integrates text information, resulting in a substantial improvement of 3.2% and 6.0% over the state-of-the-art for ‘All’ classes and ‘Old’ classes, respectively. It is worth noting that, owing to the text consistency within the same category, our text branch achieves remarkable 62.6% and 63.5% accuracy for ‘All’ classes of Imagenet-R and ImageNet-1K, respectively. The detailed results of the text branch can be seen in the Appendix.
5.3 Ablation study and analysis
Effectiveness of different components.
TABLE IV: Ablation study of different components.
TES Dual-branch CICO Stanford Cars CIFAR100
All Old New All Old New
(1) 70.0 83.4 63.5 81.1 85.0 73.3
(2) 76.2 85.3 71.7 81.0 85.3 72.3
(3) 78.5 86.8 74.5 82.1 85.5 75.5
To evaluate the effectiveness of different components, we conduct an ablation study on Stanford Cars and CIFAR100 in Tab. IV. Comparing (2) with (1), the dual-branch training strategy can effectively leverage the text features generated by TES, resulting in a 6.2% improvement on Scars’ ‘All’ classes and a 0.3% improvement on CIFAR100’s ‘Old’ classes. Furthermore, comparing (3) with (1), CICO enables the two branches to exchange information and mutually benefit from each other, resulting in remarkable improvements of 11% on Scars’ ‘New’ and 2.2% on CIFAR100’s ‘New’.
Comparison with different fusion methods.  In Tab. V, we compare our dual-branch training strategy with other methods that fuse visual and text features, including concatenation and mean. Although they may show improvements in ‘Old’ or ‘New’ classes due to the multi-modal information, we demonstrate that joint learning of the two branches is more effective as it encourages the model to have complementary and aligned discriminative capabilities for visual and text features of a same class, leading to more discriminative multi-modal prototypes and more cohesive multi-modal clusters.
TABLE V: Comparison with different fusion methods.
Dual-branch Concat Mean Stanford Cars CIFAR100
All Old New All Old New
(1) 68.9 79.1 64.0 79.9 85.5 68.7
(2) 72.0 85.0 65.6 81.1 84.3 74.8
(3) 78.5 86.8 74.5 82.1 85.5 75.5
Figure 3: Different ViT finetune strategies.
Effectiveness of text embedding synthesizer.  In order to prove that our text embedding synthesizer can generate reliable and discriminative representations, we visualize the text embeddings of CIFAR10 with t-SNE. As shown in Fig. 4, the initial text embeddings within the same class exhibit clear clustering, and the learnable embeddings further produce compacter clusters. Moreover, we introduce TES into the non-parametric GCD by straightforwardly concatenating text and image features before semi-supervised k-means classification. As in Tab. VI, with the help of text information, GCD gains about 5% average improvement on ‘All’ classes over 3 datasets, demonstrating the importance of multi-modal information in GCD task and our TES can be widely used in multiple GCD methods.
TABLE VI: Effectiveness of TES in non-parametric GCD.
FGVC-Aircraft ImageNet-100 ImageNet-R
Method All Old New All Old New All Old New
GCD-CLIP 45.3 44.4 45.8 75.8 87.3 70.0 44.3 79.0 25.8
+TES 49.6 49.3 49.8 80.0 95.1 72.4 49.4 79.4 33.5
In TES, we use a single linear layer to transform the visual embedding to pseudo tokens that serve as input to the CLIP’s text encoder. We set the number of pseudo text tokens to 7 across all datasets. Experiment results on the number of tokens and layers in TES for the CUB dataset are presented in Fig. 5, proving that a single linear layer can effectively transfer visual features into text tokens while reducing the computational cost.
Figure 5: Experiment results on the number of tokens and layers in TES for the CUB dataset.
TABLE VII: Experiments on different methods for generating pseudo text embeddings.
Method All Old New
Baseline SimGCD-CLIP 71.7 76.5 69.4
Text-Retrieval WordNet 69.8 77.1 66.2
CC3M 72.3 79.1 68.9
Text-Generation BLIP 67.1 74.3 63.5
BLIP-2 71.3 73.5 70.2
Feature-Generation TES (Ours) 77.0 78.1 76.4
TES w/o $\mathcal{L}_{align}$ 74.7 76.9 73.5
TES w/o $\mathcal{L}_{distill}$ 75.3 77.5 74.2
Our TES generates pseudo text embeddings for unlabelled data in a feature-generation approach. To demonstrate the superiority of TES, we conduct experiments that replaced the embeddings generated by TES with embeddings obtained through text-retrieval or text-generation on the CUB dataset. For text-retrieval, we retrieve the most similar text for each image in the dataset from the corpus (WordNet [45] with 117k names and CC3M [46] with 3M captions) based on the cosine similarity of the image and text embeddings. For text-generation, we use BLIP [47] and BLIP-2 [48] to perform the VQA task using the question “What’s the name of the bird in the image?”, generating text answer for each image. After obtaining the corresponding text, we encode it into text embeddings and conduct dual-branch training. As in Table VII, for fine-grained datasets like CUB, text-retrieval and text-generation methods do not enhance performance and can even be harmful. This is due to the high visual similarity in fine-grained images, leading to lower accuracy in the obtained texts. As a result, the generated text embeddings are less discriminative. Our TES, from the perspective of feature generation, optimizes through effective $\mathcal{L}_{align}$ and $\mathcal{L}_{distill}$, generating discriminative text features for the images.
5.4 Discussion about using CLIP in GCD
A key purpose of GCD is to discover novel classes, which highly rely on the initial representation discrimination provided by the backbone model. CLIP, as a large model, is trained on a huge amount of image-text pairs covering broad a set of visual concepts. Due to the strong generalization ability of CLIP, it can encode more discriminative features, making it a natural idea to introduce CLIP into the GCD task. One concern is that CLIP may have seen unknown classes/claseenames in GCD. We contend that, as a general-purpose large model, careful consideration of how to effectively unleash its potential in specific tasks is a more worthy pursuit. While CLIP may have seen unknown classes (names), the issue of being unable to use the text encoder impedes its formidable performance. Our TES provides a way to leverage the CLIP text encoder without text input, thus unlocking the multi-modal potential of CLIP.
TABLE VIII: The classnames for the NEV dataset.

### v3

@@ 4.2 Dual-branch multi-modal joint training
Since information from different modalities is exchanged and learned through CICO and injected into the visual backbone, we utilize the last-epoch visual branch for inference.
@@ 5.1 Experimental setup
CUB Stanford Cars FGVC-Aircraft CIFAR10 CIFAR100 ImageNet-100
$\mu$GCD [47] 65.7 68.0 64.6 56.5 68.1 50.9 53.8 55.4 53.0 - - - - - - - - -
LegoGCD [3] 63.8 71.9 59.8 57.3 75.7 48.4 55.0 61.5 51.7 97.1 94.3 98.5 81.8 81.4 82.5 86.3 94.5 82.1
Table 1: Results (%) on fine-grained and generic datasets. The best results are highlighted in bold.
@@
Datasets.  We evaluate our method on multiple benchmarks, including three image classification generic datasets (i.e., CIFAR 10/100 [25] and ImageNet-100 [11]), three fine-grained datasets from Semantic Shift Benchmark [46] (i.e., CUB [49], Stanford Cars [24] and FGVC-Aircraft [34]), and three challenging datasets (i.e., Herbarium 19 [44], ImageNet-R [19] and ImageNet-1K [11]). We are the first to introduce ImageNet-R into the GCD task, which contains various renditions of 200 ImageNet classes, thus challenging the GCD’s assumption that the data comes from the same domain. The data splits are reported in Supp.
Evaluation and implementation details.  Following standard evaluation protocol in [45, 51], we evaluate the performance with clustering accuracy (ACC). We use a CLIP [40] pre-trained ViT-B/16 [12] as the image and text encoder. In the first stage, we train a fully connected layer. In the second stage, we remove the projector of the image encoder, resulting in features with a dimension of 768. Other details and the pseudo-code can be found in Supp.
5.2 Comparison with state of the arts
In this section, we compare GET with several state-of-the-art methods. GCD [45] and SimGCD [51] provide paradigms for non-parametric and parametric classification, thus we replace their backbone with CLIP for a fair comparision, denoted as GCD-CLIP and SimGCD-CLIP.
Evaluation on fine-grained and generic datasets.  As shown in Tab. 1, our method achieves consistently remarkable success on all three fine-grained datasets. Specifically, we surpass SimGCD-CLIP by 5.3%, 8.5%, and 4.6% on ‘All’ classes of CUB, Stanford Cars, and Aircraft, respectively. In fine-grained datasets, the visual conceptions of distinct classes exhibit high similarity, making it challenging for classification based solely on visual information. However, their text information can provide additional discriminative information. Consequently, our GET  significantly enhances classification accuracy through the reciprocal augmentation of text and visual information flows.
Herbarium 19 ImageNet-1K ImageNet-R
Method All Old New All Old New All Old New
$k$-means [33] 13.0 12.2 13.4 - - - - - -
RS+ [18] 27.9 55.8 12.8 - - - - - -
UNO+ [14] 28.3 53.7 14.7 - - - - - -
ORCA [2] 20.9 30.9 15.5 - - - - - -
$\mu$GCD [47] 45.8 61.9 37.2 - - - - - -
LegoGCD [3] 45.1 57.4 38.4 62.4 79.5 53.8 - - -
GCD [45] 35.4 51.0 27.0 52.5 72.5 42.2 32.5 58.0 18.9
SimGCD [51] 44.0 58.0 36.4 57.1 77.3 46.9 29.5 48.6 19.4
GCD-CLIP 37.3 51.9 29.5 55.0 65.0 50.0 44.3 79.0 25.8
SimGCD-CLIP 48.9 64.7 40.3 61.0 73.1 54.9 54.9 72.8 45.3
GET (Ours) 49.7 64.5 41.7 62.4 74.0 56.6 58.1 78.8 47.0
Table 2: Results (%) on more challenging datasets.
In Tab. 1, we also present the performances for three generic datasets. Due to the low resolution of the CIFAR dataset and model biases (CLIP itself performs poorly on CIFAR100, with a zero-shot performance of 68.7), the results for novel classes are inferior compared to the DINO backbone. However, despite the inherent limitations in the discriminative capability of CLIP itself, our method still achieves an improvement of 0.4% on ‘Old’ classes of CIFAR10 and 2.2% on ‘New’ classes of CIFAR100, compared to SimGCD-CLIP. For ImageNet-100, SimGCD-CLIP has achieved an exceptionally saturated result of 90.8% on ‘All’ classes, further advancements pose considerable challenges. However, leveraging the additional modality information, GET elevates the performance ceiling to an impressive 91.7%.
Evaluation on more challenging datasets.  As shown in Tab. 2, GET outperforms all other methods for both ‘All’ and ‘New’ classes on Herb19 and ImageNet-1K datasets. In particular, our method achieves a notable improvement of 1.4% and 1.7% on ‘New’ classes of Herb19 and ImageNet-1K, respectively. Furthermore, the suboptimal performance of GCD and SimGCD with the DINO backbone on the ImageNet-R dataset highlights the difficulty of DINO in discovering new categories with multiple domains. Despite multiple domains for images of the same category, their textual information remains consistent. Our method effectively integrates text information, resulting in a substantial improvement of 3.2% and 6.0% over the sota for ‘All’ classes and ‘Old’ classes, respectively. It is worth noting that, owing to the text consistency within the same category, our text branch achieves remarkable 62.6% and 63.5% accuracy for ‘All’ classes of Imagenet-R and ImageNet-1K, respectively.
TES Dual-branch CICO Stanford Cars CIFAR100
All Old New All Old New
(1) 70.0 83.4 63.5 81.1 85.0 73.3
(2) 76.2 85.3 71.7 81.0 85.3 72.3
(3) 78.5 86.8 74.5 82.1 85.5 75.5
Table 3: Ablation study of different components.
5.3 Ablation study and analysis
Effectiveness of different components.  To evaluate the effectiveness of different components, we conduct an ablation study on SCars and CIFAR100 in Tab. 3. Comparing (2) with (1), leverage the text features generated by TES, resulting in a 6.2% improvement on SCars’ ‘All’ classes and a 0.3% improvement on CIFAR100’s ‘Old’ classes. Furthermore, comparing (3) with (1), CICO enables the two branches to exchange information and mutually benefit from each other, resulting in remarkable improvements of 11% on SCars’ ‘New’ and 2.2% on CIFAR100’s ‘New’.
Comparison with different fusion methods.  In Tab. 4, we compare our dual-branch strategy with other modality fusion methods, including concatenation and mean. Although they may show improvements due to the multi-modal information, we demonstrate that joint learning of the two branches is more effective as it encourages the model to have complementary and aligned discriminative capabilities for visual and text features of the same class, leading to more discriminative multi-modal prototypes.
Dual-branch Concat Mean Stanford Cars CIFAR100
All Old New All Old New
(1) 68.9 79.1 64.0 79.9 85.5 68.7
(2) 72.0 85.0 65.6 81.1 84.3 74.8
(3) 78.5 86.8 74.5 82.1 85.5 75.5
Table 4: Comparison with different fusion methods.
Method Total Params All Old New
Baseline SimGCD-CLIP 92.2M 71.7 76.5 69.4
Text-Retrieval WordNet 155.8M 69.8 77.1 66.2
CC3M 155.8M 72.3 79.1 68.9
VQA BLIP (ViT-L) 625.8M 67.1 74.3 63.5
BLIP-2 (opt2.7b) 3.9B 71.3 73.5 70.2
Captioning BLIP (ViT-L) 625.8M 40.5 54.6 33.4
BLIP-2 (opt2.7b) 3.9B 42.6 56.1 35.8
Feature-Generation TES (Ours) 165.1M 77.0 78.1 76.4
TES w/o $\mathcal{L}_{align}$ 165.1M 74.7 76.9 73.5
TES w/o $\mathcal{L}_{distill}$ 165.1M 75.3 77.5 74.2
Table 5: Experiments on different pseudo text embeddings.
Effectiveness of TES.  To demonstrate the superiority of TES, we conduct experiments that replaced the embeddings generated by TES with embeddings obtained through Text-Retrieval, VQA, and Captioning on the CUB dataset. For Text-Retrieval, we retrieve the most similar text for each image from two corpora (WordNet [36] and CC3M [42]) based on the cosine similarity between the image and text embeddings. We use BLIP [28] and BLIP-2 [29] to perform the Captioning and use the question “What’s the name of the bird in the image?” to perform VQA. As in Tab. 5, due to the high visual similarity in fine-grained images, category names retrieved or generated through VQA are often imprecise, making them less effective. Meanwhile, captioning methods tend to describe object poses and scenes rather than class-specific information, leading to varied captions for samples of the same class, which significantly harms category discovery. Our method achieves the best performance with moderate parameters. We provide additional experiments about TES in Supp, including its architectural design, feature distribution, and flexibility.
Results using different prompts.  In our method, we use a simple prompt: “a photo of a {CLS}." Tab. 6 presents experimental results exploring the use of alternative prompts: (1) “a photo of a {CLS}," (2) “a photo of a {CLS}, which is a type of bird/car," (3) descriptions of {CLS} generated by LLM (GPT4o-mini), and (4) average the textual features of (1) to (3). Our prompt is simple yet effective, while more finely designed prompts can further improve performance.
Prompts CUB Stanford Cars
All Old New All Old New
(1) 77.0 78.1 76.4 78.5 86.8 74.5
(2) 76.3 78.2 75.4 78.5 88.2 73.8
(3) 76.8 78.7 75.8 78.6 90.4 72.9
(4) 78.3 77.6 78.7 79.1 88.8 74.3
Table 6: Results using different prompts.
Methods NEV TV-100
All Old New All Old New
CLIP(zero-shot) 10.7 - - 1.93 - -
SimGCD 54.7 88.0 38.0 35.2 50.3 29.2
SimGCD-CLIP 79.1 96.7 70.3 55.7 75.8 47.8
GET (Ours) 85.3 96.0 80.0 57.1 77.3 49.2
Table 7: The results on the NEV and TV-100 datasets.
Discussion about using CLIP in GCD.  A key purpose of GCD is to discover novel classes, which highly rely on the initial representation discrimination provided by the backbone model. Due to the strong generalization ability of CLIP, it can encode more discriminative features, making it a natural idea to introduce CLIP into the GCD task. One concern is that CLIP may have seen unknown classes or class names in GCD. Therefore, we discuss the significance of using CLIP in GCD from three perspectives. Methodological Significance: Even CLIP is pre-trained on a vast dataset with potential overlap, its knowledge is implicit and unstructured. Effectively using it for GCD demands novel methodologies, particularly in exploring how to use the text encoder for unlabelled data. The improvements of our method over SimGCD-CLIP validate this significance. Forward-looking Significance: To evaluate the ability of category discovery methods in scenarios unseen by CLIP, we constructed a small fine-grained dataset of new energy vehicles (NEV) introduced in 2023, where CLIP lacks prior knowledge. As shown in Tab. 7, experiments on NEV and TV-100 [60](a TV series dataset that the pre-trained CLIP model has not been exposed to) demonstrate that even for categories not seen by CLIP, leveraging the text modality remains crucial for effective category discovery. This serves as a forward-looking exploration of how CLIP’s generalization can address future GCD tasks involving truly novel data. Practical Implications: Exploring CLIP’s potential in realistic scenarios is meaningful, thus, we present experiments on medical and ultra-fine-grained datasets in SSSupp. Our work lays the foundation for leveraging CLIP to address challenging GCD applications.
5.4 Qualitative results
Attention map visualization.  As in Fig. 3, compared to SimGCD-CLIP, our method additionally focuses on the feather texture of birds, which is crucial for distinguishing visually similar bird species. With the assistance of text information, the attention maps of our visual branch become more refined, focusing on more discriminative regions.
Figure 3: Attention map visualization of class tokens.
The t-SNE visualization.Fig. 4 shows the t-SNE visualization of visual and text features on the randomly sampled 20 classes of the CUB dataset. Both the visual and text features of our method exhibit clearer and compacter clusters. We provide more visualizations and cluster results in Supp.
Figure 4: The t-SNE visualizations.
6 Conclusions
In this work, we propose to leverage multi-modal information to solve the GCD task. In particular, we introduce a text embedding synthesizer to generate pseudo text embeddings for unlabelled data. Our text embedding synthesizer module makes it possible to use CLIP’s text encoder, thus unlocking the multi-modal potential for the GCD task. Meanwhile, we use a dual-branch training strategy with a cross-modal instance consistency objective, which facilitates collaborative action and mutual learning between the different modalities. Our research extending the GCD to a multi-modal paradigm and the superior performance on multiple benchmarks demonstrates the effectiveness of our method.
Limitations and future work.  A limitation of our approach is that we treat visual and text information as equally important. In fact, some samples may have richer and more discriminative visual information than textual information, and vice versa. A more appropriate approach might involve enabling the model to adaptively leverage multimodal information, autonomously assessing which modality’s information is more crucial. We will delve deeper into this aspect in our future work.
Acknowledgement
This work is funded by Shenzhen Science and Technology Program (JCYJ20240813114237048), NSFC (NO. 62206135, 62225604), Young Elite Scientists Sponsorship Program by CAST (2023QNRC001), the Fundamental Research Funds for the Central Universities (Nankai Universitiy, 070-63233085), and “Science and Technology Yongjiang 2035" key technology breakthrough plan project (2024Z120). Computation is supported by the Supercomputing Center of Nankai University.
Supplementary Material
This Supplementary Material provides the following sections:
  • Experimental setup (Sec. 8)
  • Discussion about Using CLIP in GCD (Sec. 9)
  • Additional experiments and analysis (Sec. 10)
  • Pseudo-code (Sec. 11)
  • Different pre-trained models (Sec. 12)
  • Cluster results of GET (Sec. 13)
  • Limitations and broader impact (Sec. 14)
8 Experimental setup
Datasets. We evaluate our method on multiple benchmarks, including three image classification generic datasets (i.e., CIFAR 10/100 [25] and ImageNet-100 [11]), three fine-grained datasets from Semantic Shift Benchmark [46] (i.e., CUB [49], Stanford Cars [24] and FGVC-Aircraft [34]), and three challenging datasets (i.e., Herbarium 19 [44], ImageNet-R [19] and ImageNet-1K [11]). Notice that we are the first to introduce ImageNet-R into the GCD task, which contains various renditions of 200 ImageNet classes, thus challenging the GCD’s assumption that the data comes from the same domain. For ImageNet-R, we subsample the first 100 classes as old classes, leaving the rest as new classes; the labeled dataset $\mathcal{D}_{l}$ consists of half of the old class samples, while the other half and all the new class samples are used to construct unlabelled dataset $\mathcal{D}_{u}$. Furthermore, we conduct experiments on the TV100 dataset [60], a TV series dataset that the pre-trained CLIP model has not been exposed to. We use the first 50 classes as old categories and the remaining 50 classes as new categories. As for other benchmarks, we follow the previous [45, 51] to sample $\mathcal{D}_{l}$ and $\mathcal{D}_{u}$. The details of the standard datasets we evaluate on are shown in Tab. 8.
@@ 8 Experimental setup
Table 8: The details of the standard datasets we evaluate on.
The NEV dataset  As mentioned in the main paper, we conduct a toy experiment to prove that our TES can deal with a scenario where CLIP lacks information on a specific category class. Since CLIP saw most of the visual concepts and corresponding texts before 2022, we constructed a small dataset of new energy vehicles (NEV) that appeared in 2023. As in Tab. 9, the NEV dataset contains 12 categories, each with 50 images from the Internet, and the classnames of the dataset consist of the brand and model of the car. We split them in the same way as standard benchmarks.